OpenStudy (anonymous):

In a rhombus whose side measures 36 and the smaller angle is 32°, find the length of the larger diagonal, to the nearest tenth.

OpenStudy (anonymous):

law of cosines

OpenStudy (anonymous):

yes

OpenStudy (theeric):

You don't need to know trigonometry.

OpenStudy (anonymous):

and could you show me how you get the answer

OpenStudy (anonymous):

c =\sqrt{36 ^{2}+36 ^{2}-2\left( 36 \right) \left( 36 \right) \cos \left( 180-32 \right)}\]

OpenStudy (theeric):

Here's a picture.

OpenStudy (theeric):

I misread the problem. But you should show how you got there!

OpenStudy (theeric):

Point to which diagonal is longer, @solodolo . If you were here, @pgpilot326 , I'd let you show it.

OpenStudy (theeric):

\[c =\sqrt{36 ^{2}+36 ^{2}-2\left( 36 \right) \left( 36 \right) \cos \left( 180-32 \right)}\]

OpenStudy (theeric):

I agree, but that was @pgpilot326 's equation.